ON SOME ELLIPTIC CURVES WITH LARGE SHA

HARVEY E. ROSE

Abstract. We consider a class of elliptic curves many of whose associated Shafarevich–Tate groups Sha are relatively large, and give examples of curves with o(Sha) = k^{2} for all k ≤ 100.